### v1

# PSGait: Multimodal Gait Recognition using Parsing Skeleton

Abstract
Gait recognition has emerged as a robust biometric modality due to its non-intrusive nature and resilience to occlusion. Conventional gait recognition methods typically rely on silhouettes or skeletons. Despite their success in gait recognition for controlled laboratory environments, they usually fail in real-world scenarios due to their limited information entropy for gait representations. To achieve accurate gait recognition in the wild, we propose a novel gait representation, named Parsing Skeleton. This representation innovatively introduces the skeleton-guided human parsing method to capture fine-grained body dynamics, so they have much higher information entropy to encode the shapes and dynamics of fine-grained human parts during walking. Moreover, to effectively explore the capability of the parsing skeleton representation, we propose a novel parsing skeleton-based gait recognition framework, named PSGait, which takes parsing skeletons and silhouettes as input. By fusing these two modalities, the resulting image sequences are fed into gait recognition models for enhanced individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our model. PSGait outperforms existing state-of-the-art multimodal methods. Furthermore, as a plug-and-play method, PSGait leads to a maximum improvement of 10.9% in Rank-1 accuracy across various gait recognition models. These results demonstrate the effectiveness and versatility of parsing skeletons for gait recognition in the wild, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
Index Terms:
Gait Recognition, Multimodal Fusion, Gait Representation, Human Parsing
I Introduction
Gait recognition aims to identify the target person by the motion pattern of human walking. Compared with face and fingerprint, gait has the advantages of being remotely accessible, non-contacting, and hard to disguise, which makes it a promising biometric for applications like ID authentication and social security[1, 2].
Existing gait representations for gait recognition primarily rely on silhouette sequences [3], Gait Energy Images (GEIs)[4], 2D skeletons[5], 3D skeletons[6], etc. Although these representations perform well in laboratory environments, they often fail in unconstrained situations, i.e., gait recognition in the wild [7, 8]. This failure is attributed to the low information entropy of binary silhouettes or 2D/3D skeletons, which are insufficient for effectively encoding the shapes and dynamics of the human body during walking. Recent research has increasingly focused on multimodal fusion methods in gait recognition[9, 10, 8, 11, 12], combining silhouettes and skeletons to create a more comprehensive representation of human gait. For example, Skeletongait++[9] utilized the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism. Although this approach has shown improvements, it relied on Gaussian points to represent the skeleton, neglecting detailed information about the body parts between these points. In gait recognition, the motion dynamics of different body parts vary significantly, and this disparity provides valuable complementary information for distinguishing individuals. Therefore, it is crucial to explore new representations with higher information entropy for gait recognition in the wild. Recent studies have demonstrated that human parsing provides a fine-grained and more comprehensive representation of gait, significantly enhancing the performance of gait recognition models[13, 14, 15, 16]. Although human parsing has shown promising results in gait recognition, conventional methods typically rely on RGB images, which are highly sensitive to factors like clothing variations and environmental occlusion, thereby reducing the accuracy of gait recognition in the wild.
Figure 1: The comparison of various gait representations. Only a single frame is displayed for brevity. Result shows the average Rank-1 on CCPG[17] for different representation inputs to GaitBase[18].
To address these issues, we propose a novel gait representation named Parsing Skeleton, which can be easily extracted through image processing methods based on the skeleton. Compared to existing gait recognition representations, the parsing skeleton offers three key advantages: 1) it contains higher information entropy at the pixel level, encoding more details about human motion; 2) it represents part-level knowledge, enabling models to learn fine-grained dynamics of different body parts during walking; 3) being derived from skeleton data, it improves the robustness and accuracy of gait recognition in real-world scenarios, outperforming existing RGB-based parsing methods. To effectively explore the capability of the parsing skeleton, we propose a parsing skeleton-based gait recognition framework named PSGait, which takes parsing skeletons and silhouettes as input. By fusing these two modalities through overlaying, the resulting image sequences are fed into gait recognition models for individual differentiation. We conduct comprehensive benchmarks on various datasets to evaluate our method against existing gait recognition approaches. PSGait achieves a maximum improvement of 5.2% in Rank-1 accuracy compared to the state-of-the-art multimodal method. Furthermore, replacing silhouettes with the fusion of silhouettes and parsing skeletons maximally leads to a 10.9% improvement in Rank-1 accuracy across various gait recognition models, underscoring the significant contribution of the plug-and-play parsing skeleton to gait recognition in real-world scenarios. These results validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
Figure 2: Overview of the PSGait framework. During data preprocessing, the silhouettes and skeletons are generated from video frames. In multimodal fusion, after the parsing section, the parsing skeletons are fused with silhouettes. The fused data is then fed into the gait recognition model for individual differentiation. The parsing section refines body part representations through selected points, creating circular and linear representations to enhance the robustness and accuracy of gait recognition in the wild. Both the data preprocessing and gait recognition models are interchangeable.
In summary, the contributions of this paper are as follows:
  • To the best of our knowledge, this is the first work to explore a simple yet effective gait representation for gait recognition, namely Parsing Skeleton, which demonstrates higher information entropy compared to existing silhouette- or skeleton-based representations.
  • We introduce PSGait, a novel gait recognition framework based on the parsing skeleton, which extracts comprehensive gait features with parsing skeletons and silhouettes, thereby enhancing the feature robustness in the wild.
  • Extensive experiments across various datasets validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art method for multimodal gait recognition.
II Related Work
Multimodal gait recognition aims to fuse multiple modalities to create a more comprehensive representation of gait. Specifically, SMPLGait[8] leveraged 3D geometric information from the SMPL model to enhance the gait appearance feature learning. BiFusion[10] integrated silhouettes and skeletons, using a graph convolutional network (GCN) with a multi-scale gait graph (MSGG) to extract complementary features from both modalities. Skeletongait++[9] employed the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
Human parsing in gait recognition typically relies on representations that contain fine-grained semantic information about individual body parts. These methods achieve simple but efficient human body parsing based on RGB images. Specifically, ParsingGait[13] combined CNNs and GCNs to effectively capture both global and local gait features, thereby enhancing gait recognition performance in real-world scenarios. GaitParsing[14] decomposed gait into specific body parts using human semantic parsing and employed a dual-branch feature extraction network to process both holistic and part-level features. LandmarkGait[16] introduced an unsupervised method that converts silhouettes into consistent landmarks for fine-grained parsing, utilizing a multi-scale feature extractor to capture both global and part-level features for gait recognition.
III Method
The PSGait proposed in this study is illustrated in Fig. 2. First, during data preprocessing, the silhouette and skeleton for each frame of the gait video sequence are obtained through segmentation and pose estimation. Next, after the parsing section, the Parsing Skeletons are fused with silhouettes to obtain the multimodal gait representation. Finally, this multimodal gait representation, which encapsulates comprehensive information, is fed into the gait recognition model to enhance its classification performance.
III-A Parsing Section
The main objective of this section is to obtain representations with fine-grained semantic information of various body parts from the skeleton. First, the point selection module involves two steps: data cleaning and human body part selection. To mitigate the impact of erroneous results from the pose estimation during data preprocessing, a confidence threshold $\tau$ is set to filter out unreliable points. Only points with confidence $c_{i}$ exceeding the threshold $\tau$ are used. Additionally, any points outside the valid region ($H\times W$) are disregarded. Therefore, a valid point ($x_{i},y_{i}$) must satisfy the following condition:
[…]
According to the joint point configuration in COCO17, we selected the corresponding joint points for each body part, including the head and limbs. For the head, three points were selected: the nose and both ears. For each limb, the joint points at both ends were chosen. This results in the following mapping between human body joints $k_{i}$ and their corresponding body parts $P_{j}$:
$h(P_{j})=\{k_{1},k_{2},\dots,k_{m}\}$ (2)
When rendering, we need to ensure that the parsing skeleton’s information is effectively captured by the CNN when it is fed into the gait recognition model with uniform size. Additionally, different colors $c$ are assigned to each body part to enhance the fine-grained information of the human body. For the head representation, we introduce circles centered at the three key head points ($x,y$), with a specified radius $r$, encompassing the around the pixel $(u,v)$:
$\text{circle}(x,y,r,c)=\{(u,v)\mid(u-x)^{2}+(v-y)^{2}\leq r^{2}\}$ (3)
For the representation of each limb, a line of specified width is parsed on the image:
$\text{line}((x_{1},y_{1}),(x_{2},y_{2}),w,c)=\left\{(u,v)\mid d\leq\frac{w}{2}\right\}$ (4)
where $d$ represents the distance of the pixel $(u,v)$ from the line, $c$ denotes the color assigned to this part, and $w$ is the width of this line.
III-B Fusion Strategy
In this study, we investigate two distinct fusion strategies for combining parsing skeleton and silhouette in representation learning. The objective of the fusion is to generate an effective input that captures both the global human shape from the silhouette and the fine-grained local structural relationships from the parsing skeleton.
Overlay-Based Fusion: The first strategy is a straightforward overlay operation, where the parsing skeleton representation (lines and circles) is directly superimposed onto the silhouette image $M(x,y)$. Specifically, the fused image $F_{\text{overlay}}(x,y)$ is defined as follows:
$F_{\text{overlay}}(x,y)=\max\left(M(x,y),S(x,y)\right),$ (5)
where $M(x,y)$ is the binary silhouette image, and $S(x,y)$ represents the parsing skeleton. The $\max$ operator ensures that the parsing skeleton structure is not overwritten by the silhouette but is integrated naturally. The overlay method produces an RGB fused image, where the silhouette and parsing skeleton share the same spatial space, allowing the network to jointly learn the global shape features and structural details with minimal processing complexity.
Multi-Channel Fusion: In the second strategy, we introduce a more granular representation by separating each parsing skeleton component (lines and circles) into independent channels. Specifically, each limb and head point are encoded in separate channels, and the silhouette $M(x,y)$ is placed in the final channel. This results in a multi-channel representation $F_{\text{multi}}$, where the fused feature map has 12 channels:
$F_{\text{multi}}(x,y,c)=\begin{cases}S_{c}(x,y),&\text{if }c=1,2,\dots,11\\ M(x,y),&\text{if }c=12\end{cases}$ (6)
where $S_{c}(x,y)$ represents the $c$-th parsing skeleton component (either a line or a circle), and the 12th channel contains the silhouette image. While this strategy provides a more explicit and structured representation, enabling the network to focus on individual components of the parsing skeleton and silhouette, it inherently increases the complexity of the input and may dilute spatial relationships across channels.
III-C Align Center
After the data fusion stage, the center alignment operation is performed to ensure that the human body contour and the resolved skeleton are centered in the image under the premise of uniform input image size. This alignment facilitates subsequent model processing and feature extraction. Given a binary silhouette image $M(x,y)$, where $M(x,y)\in\{0,1\}$, the upper bound $y_{\text{top}}$ and lower bound $y_{\text{btm}}$ of the human body region are calculated as follows:
$y_{\text{top}}=\min\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (7)
$y_{\text{btm}}=\max\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (8)
Next, the image is cropped based on the upper and lower bounds of the human region and resized to the target height while preserving the aspect ratio, resulting in the image $M^{\prime}(x,y)$. The center of the horizontal axis is calculated as follows:
$x_{c}=\frac{\sum_{x,y}x\cdot M^{\prime}(x,y)}{\sum_{x,y}M^{\prime}(x,y)}$ (9)
Using the horizontal axis center $x_{c}$, the image is cropped to a target size of $H^{\prime}\times W^{\prime}$.
TABLE I: Quantitative comparison of gait recognition methods on the SUSTech1K[19] is presented. The best performances are indicated in bold, and the second-best methods are underlined. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Probe Sequence(R-1) Overall
Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
Skeleton GaitGraph2(CVPRW2022)[20] 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
GaitTR(ES2023)[21] 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
Silhouette GaitSet(AAAI2019)[22] 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
GaitPart(CVPR2019)[23] 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
GaitGL(ICCV2021)[24] 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
GaitBase(CVPR2023)[18] 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
DeepGaitV2(Arxiv2023)[25] 83.5 79.5 46.3 76.8 79.1 78.5 81.1 27.3 77.4 90.2
[…]
Skeleton+Silhouette BiFusion(MTA2023)[10] 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
SkeletonGait++†(AAAI2024)[9] 84.7 82.4 46.7 81.8 81.2 82.6 85.1 44.3 81.2 93.2
[…]
TABLE II: Quantitative comparison of gait recognition methods is conducted across two authoritative datasets, CCPG[17] and Gait3D[8]. †We modified Skeletongait++ to match the settings used in GaitBase.
Input Method Testing Datasets
CCPG Gait3D
CL UP DN BG Mean Rank-1 Rank-5 mAP mINP
Skeleton GaitGraph2(CVPRW2022)[20] 5.0 5.3 5.8 6.2 5.6 11.1 - - -
GaitTR(ES2023)[21] 15.7 18.3 18.5 17.5 17.5 6.6 - - -
Silhouette GaitSet(AAAI2019)[22] 60.2 65.2 65.1 68.5 64.8 36.7 58.3 30.0 17.3
GaitPart(CVPR2020)[23] 64.3 67.8 68.6 71.7 68.1 28.2 47.6 21.6 12.4
GaitBase(CVPR2023)[18] 71.6 75.0 76.8 78.6 75.5 64.6 - - -
[…]
Skeleton+Silhouette SkeletonGait++†(AAAI2024)[9] 77.0 81.4 83.0 88.8 82.6 64.8 80.5 54.8 36.1
[…]
IV Experience
IV-A Datasets
We evaluated our proposed method on three commonly used datasets: SUSTech1K[19] and CCPG[17], which are collected in fully constrained laboratory environments, and Gait3D[8], which is collected from real-world scenarios.
SUSTech1K[19] contains gait data from 1,050 subjects, covering 8 different conditions: normal (NM), bag (BG), clothes changing (CL), carrying (CR), umbrella (UB), uniform (UF), occlusion (OC), and night (NT), as well as 12 viewpoints (ranging from 0° to 315°). The diversity and granularity of the dataset make SUSTech1K[19] an ideal platform for studying the performance of gait recognition in complex environments.
CCPG[17] was designed to address the challenge of gait recognition under clothing change conditions and is particularly suitable for dynamic person recognition in video surveillance. The dataset includes 200 identities and more than 16,000 sequences, with each identity captured from 10 viewpoints under seven different clothing change conditions, covering both indoor and outdoor environments. It provides two data modalities: RGB images and silhouette images.
Gait3D[8] is a large-scale gait dataset collected in real-world environments. It was recorded in a supermarket and contains 1,090 hours of video footage with a resolution of 1,920×1,080 and a frame rate of 25 FPS. Gait3D provides four modalities: silhouettes, 2D and 3D skeletons, and 3D meshes.
TABLE III: Impact of circle radius and line width on the SUSTech1K[19] dataset using GaitBase[18].
Circle Radius Line Width Probe Sequence(R-1) Overall
Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
3 3 82.1 80.2 47.0 77.9 80.8 79.4 87.1 33.8 78.8 91.7
10 12 86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
20 24 84.6 83.1 51.3 80.8 83.0 81.2 90.0 43.1 81.6 93.7
TABLE IV: Impact of input modalities and fusion strategies on the SUSTech1K[19] dataset using GaitBase[18].
Skeleton Silehouette Overlay Probe Sequence(R-1) Overall
Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
52.5 50.1 35.2 48.4 40.3 51.3 61.1 22.0 50.3 72.5
81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
82.7 81.2 47.8 78.9 81.3 79.8 88.2 35.1 79.7 92.2
86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
IV-B Implementation Details
For data preprocessing, the original data is used to extract silhouette information from all three datasets. Since skeleton data is not available for the CCPG dataset, we utilize the Sapiens[26] method for pose estimation. The original skeleton data is used for the other two datasets, and the format of points follows COCO17. A confidence threshold of 0.1 is applied.
In the parsing stage, to ensure that the Parsing Skeleton features of the processed image can be effectively captured by the CNN, the radius of the circle representing the head and the width of the line segment representing the limbs are determined based on the image size ratio in each dataset. Specifically, for the SUSTech1K[19] dataset, a radius of 10 and a line width of 12 are used. For the CCPG[17] dataset, the radius and line width are set to 16 and 20, respectively. Similarly, for the Gait3D[8] dataset, the radius and line width are also set to 16 and 20, respectively. The colors selected for parsing each body part are shown in Fig. 2.
To evaluate the versatility of our method, we apply it to two state-of-the-art gait recognition models: GaitBase[18] and DeepGaitV2[25]. The size of the fused image is adjusted to $64\times 44$ before being input into the gait recognition model. For model training, we largely follow the parameter settings from OpenGait[18], an excellent open-source framework in the field of gait recognition. Specifically, for the SUSTech1K[19] dataset, the batch size is set to $(8,8)$, and learning rate milestones are scheduled at 20k, 30k, and 40k steps, with a total of 50k training steps. For the CCPG[17] dataset, we use a batch size of $(8,16)$ with learning rate milestones at 20k, 40k, and 50k steps, and a total of 60k iterations. For the Gait3D[8] dataset, the batch size is set to $(32,4)$, with learning rate milestones occurring at 20k, 40k, and 50k steps, and a total of 60k iterations. The optimization process is configured using the SGD solver. Specifically, the learning rate $lr$ is initialized at 0.1, providing an appropriate step size for updating the model weights during training. A momentum value of 0.9 is employed to accelerate convergence by reducing oscillations and incorporating a weighted history of previous gradients.
IV-C Comparison with State-of-the-art Methods
As shown in Table I and Table II, we evaluate the performance on three datasets. The results highlight the strong competitiveness and versatility of our method.
Competitiveness. We compare and validate the advantages of our method across three datasets. On the SUSTech1K[19] dataset (Table I), compared to SkeletonGait++, Ours+GaitBase improves the Rank-1 accuracy from 81.2% to 83.4% (+2.2%). This demonstrates the effectiveness of our method under the same baseline settings. Furthermore, Ours+DeepGaitV2 achieves the best overall performance on SUSTech1K, attaining a Rank-1 accuracy of 87.9% and a Rank-5 accuracy of 94.8%. On the CCPG[17] dataset (Table II), compared to SkeletonGait++[9], Ours+GaitBase improves the mean accuracy from 82.6% to 86.4% (+3.8%). On the Gait3D[8] dataset (Table II), Rank-1 accuracy is improved from 64.8% to 70.0% (+5.2%). Additionally, Ours+GaitBase and Ours+DeepGaitV2 achieve the best overall performance on CCPG and Gait3D, with a mean accuracy of 86.4% on CCPG and a Rank-1 accuracy of 85.6% and mAP of 69.4% on Gait3D, outperforming all other state-of-the-art methods. These improvements underscore the robustness and adaptability of our approach across diverse benchmarks and evaluation metrics.
Versatility. We further demonstrate the versatility of our approach by integrating it into two baseline models, GaitBase[18] and DeepGaitV2[25], and comparing the resulting performance improvements. On the three datasets mentioned above, our method provides a significant boost over the existing baselines. Specifically, on the SUSTech1K dataset[19], Ours+GaitBase improves the Rank-1 accuracy from 76.1% to 83.4% (+7.3%), while Ours+DeepGaitV2 enhances the Rank-1 accuracy from 77.4% to 84.6% (+7.2%). On the CCPG[17] dataset, integrating our method with GaitBase increases the mean accuracy from 75.5% to 86.4% (+10.9%). For DeepGaitV2, the mean accuracy improves from 83.3% to 85.6% (+2.3%). On Gait3D[8], Ours+GaitBase improves the Rank-1 accuracy from 64.6% to 70.0% (+5.4%). For DeepGaitV2, the Rank-1 accuracy increases from 74.4% to 76.2% (+1.8%), and the mAP rises from 65.8% to 69.4% (+3.6%). These consistent performance improvements across multiple datasets and baseline models highlight the strong versatility of our method. It effectively enhances the performance of any gait recognition model, making it plug-and-play and universally applicable to various scenarios and datasets.
IV-D Ablation Study
To validate the effectiveness of our method, we conducted ablation studies on the SUSTech1K[19] dataset using GaitBase[18] as the baseline gait recognition model. We conducted experiments to assess the impact of key parameters, including circle radius, line width, input modalities, and different data fusion strategies.
Influence of Circle Radius and Line Width. As shown in Table III, we evaluated the impact of circle radius and line width on recognition performance. A smaller setting results in a relatively low Rank-1 accuracy of 82.1%, while a moderate setting achieves the best Rank-1 accuracy of 86.1% and Rank-5 accuracy of 94.3%. However, further increasing the radius and width, slightly reduces Rank-1 accuracy to 84.6%, indicating that finding an optimal balance is crucial for performance. The results indicate that choosing an appropriate circle radius and line width is critical for effectively extracting parsing skeleton features. If these values are too small, the parsing skeleton cannot be effectively captured by the CNN after resizing the image. Conversely, if they are too large, the parsing skeleton may obscure the silhouette, impeding the extraction of human shape features by gait recognition models.
Impact of Data Fusion Strategies. We further analyzed the contribution of different modalities and fusion methods (parsing skeleton, silhouette, and overlay) to overall performance. As shown in Table IV, using parsing skeletons or silhouettes alone does not yield particularly high performance, indicating that relying on either type of data alone is insufficient for accurate gait recognition. Incorporating silhouette information significantly improves performance, with Rank-1 increasing to 79.7% and Rank-5 reaching 92.2%. In this case, the data fusion strategy employed is a multi-channel fusion method. When parsing skeleton and silhouette are overlaid, the model achieves the best performance, with a Rank-1 accuracy of 83.4% and a Rank-5 accuracy of 94.3%. The overlay fusion strategy capitalizes on the complementary strengths of skeleton and silhouette data, allowing the model to capture both global shape and local structural details. These results demonstrate that our fusion strategy is both effective and achieves an optimal balance between accuracy and robustness.
V Conclusion
In this paper, we propose a novel gait representation, Parsing Skeleton, which leverages skeleton-guided human parsing to capture fine-grained body dynamics. This representation addresses the limitations of traditional gait recognition methods that rely on silhouettes or skeletons with low information entropy, by encoding more detailed and comprehensive features of human motion. To effectively exploit the parsing skeleton, we propose PSGait, a new gait recognition framework that integrates parsing skeletons with silhouettes to enhance feature robustness and improve individual differentiation, particularly in unconstrained, real-world scenarios. Extensive experiments across various datasets demonstrate that PSGait outperforms existing state-of-the-art multimodal gait recognition methods. As a plug-and-play method, PSGait leads to significant performance improvements across various models. This work validates the effectiveness and versatility of the parsing skeleton representation for real-world gait recognition tasks, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.

### v2

@@
-PSGait: Multimodal Gait Recognition using Parsing Skeleton
-Abstract
+PSGait: Gait Recognition using Parsing Skeleton
+Abstract.
 Gait recognition has emerged as a robust biometric modality due to its non-intrusive nature and resilience to occlusion.
@@
 We conduct comprehensive benchmarks on various datasets to evaluate our model.
-PSGait outperforms existing state-of-the-art multimodal methods.
+PSGait outperforms existing state-of-the-art multimodal methods that utilize both skeleton and silhouette inputs while significantly reducing computational resources.
 Furthermore, as a plug-and-play method, PSGait leads to a maximum improvement of 10.9% in Rank-1 accuracy across various gait recognition models.
-These results demonstrate the effectiveness and versatility of parsing skeletons for gait recognition in the wild, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
-Index Terms:
-Gait Recognition, Multimodal Fusion, Gait Representation, Human Parsing
-I Introduction
-Gait recognition aims to identify the target person by the motion pattern of human walking.
-Compared with face and fingerprint, gait has the advantages of being remotely accessible, non-contacting, and hard to disguise, which makes it a promising biometric for applications like ID authentication and social security[1, 2].
-Existing gait representations for gait recognition primarily rely on silhouette sequences [3], Gait Energy Images (GEIs)[4], 2D skeletons[5], 3D skeletons[6], etc.
-Although these representations perform well in laboratory environments, they often fail in unconstrained situations, i.e., gait recognition in the wild [7, 8].
-This failure is attributed to the low information entropy of binary silhouettes or 2D/3D skeletons, which are insufficient for effectively encoding the shapes and dynamics of the human body during walking.
-Recent research has increasingly focused on multimodal fusion methods in gait recognition[9, 10, 8, 11, 12], combining silhouettes and skeletons to create a more comprehensive representation of human gait.
-For example, Skeletongait++[9] utilized the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
-Although this approach has shown improvements, it relied on Gaussian points to represent the skeleton, neglecting detailed information about the body parts between these points.
-In gait recognition, the motion dynamics of different body parts vary significantly, and this disparity provides valuable complementary information for distinguishing individuals.
-Therefore, it is crucial to explore new representations with higher information entropy for gait recognition in the wild.
-Recent studies have demonstrated that human parsing provides a fine-grained and more comprehensive representation of gait, significantly enhancing the performance of gait recognition models[13, 14, 15, 16].
-Although human parsing has shown promising results in gait recognition, conventional methods typically rely on RGB images, which are highly sensitive to factors like clothing variations and environmental occlusion, thereby reducing the accuracy of gait recognition in the wild.
-Figure 1: The comparison of various gait representations. Only a single frame is displayed for brevity. Result shows the average Rank-1 on CCPG[17] for different representation inputs to GaitBase[18].
-To address these issues, we propose a novel gait representation named Parsing Skeleton, which can be easily extracted through image processing methods based on the skeleton.
-Compared to existing gait recognition representations, the parsing skeleton offers three key advantages: 1) it contains higher information entropy at the pixel level, encoding more details about human motion; 2) it represents part-level knowledge, enabling models to learn fine-grained dynamics of different body parts during walking; 3) being derived from skeleton data, it improves the robustness and accuracy of gait recognition in real-world scenarios, outperforming existing RGB-based parsing methods.
-To effectively explore the capability of the parsing skeleton, we propose a parsing skeleton-based gait recognition framework named PSGait, which takes parsing skeletons and silhouettes as input.
-By fusing these two modalities through overlaying, the resulting image sequences are fed into gait recognition models for individual differentiation.
-We conduct comprehensive benchmarks on various datasets to evaluate our method against existing gait recognition approaches.
-PSGait achieves a maximum improvement of 5.2% in Rank-1 accuracy compared to the state-of-the-art multimodal method.
-Furthermore, replacing silhouettes with the fusion of silhouettes and parsing skeletons maximally leads to a 10.9% improvement in Rank-1 accuracy across various gait recognition models, underscoring the significant contribution of the plug-and-play parsing skeleton to gait recognition in real-world scenarios.
-These results validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
-Figure 2: Overview of the PSGait framework.
+These results demonstrate that Parsing Skeleton offers a lightweight, effective, and highly generalizable representation for gait recognition in the wild.
+Gait Recognition, Gait Representation, Biometric Modality, Parsing Skeleton, Human Parsing, Multimodal Gait Recognition
+copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 33st ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.ccs: Computing methodologies Biometrics
+1. INTRODUCTION
+Figure 1.
+The comparison of various gait representations.
+Only a single frame is displayed for brevity.
+Result shows the average Rank-1 on CCPG(Li et al., 2023b) for different representation inputs to GaitBase(Fan et al., 2023b).
+Gait recognition aims to identify individuals based on their walking patterns.
+Compared to other biometric modalities such as face and fingerprint, gait is non-intrusive, contactless, and can be captured at a distance without subject cooperation.
+These advantages make gait particularly appealing for real-world applications, including person identification, access control, and public security surveillance in unconstrained environments (Niyogi and Adelson, 1994; Wan et al., 2018; Wang et al., 2003; Wu et al., 2016; Sarkar et al., 2005; Nixon et al., 1996, 2010).
+Existing gait recognition methods typically rely on representations such as silhouette sequences (Yu et al., 2006), Gait Energy Images (GEIs) (Han and Bhanu, 2005), 2D skeletons (Teepe et al., 2021), or 3D skeletons (Liao et al., 2020).
+Among them, silhouette-based approaches are widely adopted due to their simplicity and effectiveness in controlled environments.
+However, as binary images, silhouettes offer limited information entropy—capturing only coarse body contours while discarding fine-grained motion and structural details.
+They are also highly sensitive to clothing variations, occlusion, and background noise, which often results in incomplete or distorted inputs.
+Furthermore, silhouette quality heavily relies on background subtraction algorithms, making them vulnerable to domain shifts and environmental inconsistencies across datasets.
+The lack of semantic body part information further restricts the model’s ability to learn localized motion patterns, ultimately limiting the robustness and discriminability of silhouette-based representations in the wild(Zheng et al., 2023).
+Skeleton-based representations provide an alternative that abstracts human motion through joint coordinates.
+Recent studies have explored skeleton-based gait recognition using Graph Convolutional Networks (GCNs)(Teepe et al., 2021, 2022; Zhang et al., 2023; Fu et al., 2023), but performance still lags behind image-based methods.
+We argue this gap stems from the inherently low-dimensional and sparse nature of coordinate-based skeletons (e.g., 17 joints in COCO format), which results in low information entropy.
+Such sparse representations lack the spatial density and structural detail needed to model complex body dynamics during walking.
+Additionally, GCN-based methods are often constrained by limited receptive fields and struggle to capture long-range spatial dependencies, making it challenging to extract both global posture and local motion cues effectively.
+To mitigate the limitations of single-modality representations, recent works have increasingly focused on multimodal fusion strategies that integrate silhouettes and skeletons (Fan et al., 2024; Peng et al., 2024; Zheng et al., 2022; Cui and Kang, 2023; Li et al., 2023a).
+For instance, SkeletonGait++ (Fan et al., 2024) proposed a Skeleton Map to visually encode joint coordinates and fused it with silhouettes via attention mechanisms.
+While this approach improves representation capacity, it still uses Gaussian points to encode joints and overlooks the motion dynamics of body regions between these points.
+Given that different body parts exhibit distinct motion trajectories, capturing part-level movement is crucial for discriminative gait recognition.
+This motivates the need for richer representations with higher information entropy that can better characterize human motion in unconstrained settings.
+Human parsing has recently been explored as a promising avenue to enhance gait representation through fine-grained semantic segmentation (Zheng et al., 2023; Wang et al., 2023a; Zheng et al., 2024; Wang et al., 2023b).
+Parsing-based methods can provide part-level spatial detail and capture diverse motion patterns across different body regions.
+However, most existing approaches rely on RGB images, which are inherently sensitive to illumination changes, clothing variations, and occlusion—factors that are prevalent in real-world scenarios.
+This undermines the reliability and robustness of parsing-based representations in the wild.
+To address these challenges, we propose a novel gait representation called Parsing Skeleton, which introduces skeleton-guided human parsing to generate dense, fine-grained body-part maps.
+Parsing Skeleton offers three key advantages over existing representations: (1) it exhibits higher pixel-level information entropy by encoding detailed motion cues; (2) it explicitly captures part-level semantic knowledge, enabling the model to distinguish diverse motion dynamics of different body regions; and (3) it inherits the robustness of skeletons to environmental noise and domain shifts, making it more reliable in real-world scenarios than RGB-based parsing methods.
+Unlike sparse coordinate-based skeletons, the Parsing Skeleton transforms joint positions into structured visual patterns, representing body parts (e.g., limbs and head) as colored line segments and circles.
+This enhances spatial density, preserving both local motion cues and global context.
+Its image-based nature makes it highly compatible with CNNs, which excel at learning spatially invariant features.
+Additionally, the Parsing Skeleton encodes structural priors and can be easily fused with silhouettes.
+In contrast, GCNs struggle with long-range dependencies and lack spatial expressiveness, making CNNs a more effective backbone.
+To explore the potential of the Parsing Skeleton, we propose PSGait, a gait recognition framework that integrates Parsing Skeletons and silhouettes.
+By fusing these two modalities, the resulting image sequences are fed into gait recognition models for individual differentiation.
+We conduct comprehensive benchmarks on several datasets, demonstrating that PSGait achieves a maximum improvement of 5.2% in Rank-1 accuracy compared to the state-of-the-art multimodal methods that utilize both skeleton and silhouette inputs while significantly reducing computational resources.
+Moreover, replacing silhouettes with the fusion of silhouettes and Parsing Skeletons maximally leads to a 10.9% improvement in Rank-1 accuracy across various models, underscoring the significant contribution of the Parsing Skeleton to gait recognition in real-world scenarios.
+In summary, the contributions of this paper are as follows:
+  • We propose Parsing Skeleton, a novel and lightweight gait representation that offers higher information entropy and fine-grained body-part awareness compared to traditional silhouette- and skeleton-based representations.
+  • We introduce PSGait, a novel gait recognition framework based on the Parsing Skeleton, which extracts comprehensive gait features with Parsing Skeletons and silhouettes, thereby enhancing the feature robustness in the wild.
+  • Extensive experiments on multiple datasets validate the lightweight, effectiveness and generalization nature of Parsing Skeleton.
+PSGait achieves state-of-the-art performance among multimodal gait recognition methods while maintaining high computational efficiency.
+2. RELATED WORK
+In this section, we first introduce the existing gait representations. Then, the gait recognition methods are briefly surveyed. Finally, we summarize the human parsing methods in gait recognition.
+2.1. Gait Representations
+Gait recognition is a rapidly evolving field, with a primary focus on extracting distinctive gait features from visual inputs.
+The majority of gait representations are derived from traditional RGB images, including raw RGB images, binary silhouettes, optical flow images, 2D/3D skeletons, and human meshes.
+These representations are often susceptible to extraneous noise, stemming from factors such as background clutter, clothing variations, and external accessories (e.g., bags).
+To mitigate such influences, preprocessing stages such as background subtraction, silhouette extraction, and normalization are typically employed.
+Alternatively, end-to-end learning approaches, which directly learn robust gait features from raw data, are also utilized to overcome such challenges.
+Beyond RGB images, recent studies have explored alternative sensor modalities, such as LiDAR (Shen et al., 2023) and event cameras (Wang et al., 2022), which capture temporal and spatial information with high precision.
+However, the integration of these sensors remains limited due to their relatively low adoption in existing CCTV systems, thereby rendering them unsuitable for large-scale deployment in surveillance applications at present.
+2.2. Gait Recognition Methods
+According to the classical taxonomy, gait recognition methods can be broadly classified into two categories: model-based and appearance-based methods.
+Model-based gait recognition methods utilize the underlying structure of the human body as input, such as the estimated 2D / 3D skeleton and human mesh.(Wu et al., 2025; Ariyanto and Nixon, 2011; Yam et al., 2004; Urtasun and Fua, 2004; Yamauchi et al., 2009; Zhao et al., 2006) With extremely excluding visual clues, these gait representations, which are formally parameterized as coordinates of human joints or customized vectors in most cases, are theoretically ‘clean’ against factors like carrying and dressing items.
+In recent literature, PoseGait (Liao et al., 2020) combines the 3D skeleton data with hand-crafted characteristics to overcome the viewpoint and clothing variations, GaitGraph (Teepe et al., 2021) introduces a graph convolution network for 2D skeleton-based gait representation learning, HMRGait (Li et al., 2020) fine-tunes a pre-trained human mesh recovery network to construct an end-to-end SMPL-based model.
+Despite the advances achieved on indoor OU-MVLP(Takemura et al., 2018), previous model-based methods still have not exhibited competitive performance on real-world gait datasets.
+Appearance-based gait recognition methods mostly learn gait features from silhouette or RGB images, leveraging informative visual characteristics.(Lin et al., 2021; Zhang et al., 2021, 2019; Han and Bhanu, 2005; Shiraga et al., 2016; Wu et al., 2016; Huang et al., 2021) With the advent of deep learning, current appearance-based approaches primarily concentrate on spatial feature extraction and gait temporal modeling.
+Specifically, GaitSet (Chao et al., 2021) innovatively treats the gait sequence as a set and employs a maximum function to compress the sequence of frame-level spatial features.
+Due to its simplicity and effectiveness, GaitSet has emerged as one of the most influential gait recognition works in recent years.
+GaitPart (Fan et al., 2020) meticulously explores the local details of input silhouettes and models temporal dependencies using the Micro-motion Capture Module.
+GaitGL (Lin et al., 2022) argues that spatially global gait representations often overlook important details, while local region-based descriptors fail to capture relationships among neighboring parts.
+Consequently, GaitGL introduces global and local convolution layers.
+More recently, DeepGaitV2 (Fan et al., 2023a) presents a unified perspective to explore how to construct deep models for outdoor gait recognition, bringing a breakthrough improvement on the challenging Gait3D and GREW.
+Additionally, there are also some progressive multi-modal gait recognition methods aiming to fuse multiple modalities to create a more comprehensive representation of gait.
+Specifically, SMPLGait (Zheng et al., 2022) exploited the 3D geometrical information from the SMPL model to enhance the gait appearance feature learning.
+BiFusion(Peng et al., 2024) integrated silhouettes and skeletons, using a graph convolutional network (GCN) with a multi-scale gait graph (MSGG) to extract complementary features from both modalities.
+Skeletongait++(Fan et al., 2024) employed the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
+Figure 2.
+Overview of the PSGait framework.
 During data preprocessing, the silhouettes and skeletons are generated from video frames.
@@
 Both the data preprocessing and gait recognition models are interchangeable.
-In summary, the contributions of this paper are as follows:
-  • To the best of our knowledge, this is the first work to explore a simple yet effective gait representation for gait recognition, namely Parsing Skeleton, which demonstrates higher information entropy compared to existing silhouette- or skeleton-based representations.
-  • We introduce PSGait, a novel gait recognition framework based on the parsing skeleton, which extracts comprehensive gait features with parsing skeletons and silhouettes, thereby enhancing the feature robustness in the wild.
-  • Extensive experiments across various datasets validate the effectiveness and versatility of the parsing skeleton, establishing PSGait as a new state-of-the-art method for multimodal gait recognition.
-II Related Work
-Multimodal gait recognition aims to fuse multiple modalities to create a more comprehensive representation of gait.
-Specifically, SMPLGait[8] leveraged 3D geometric information from the SMPL model to enhance the gait appearance feature learning.
-BiFusion[10] integrated silhouettes and skeletons, using a graph convolutional network (GCN) with a multi-scale gait graph (MSGG) to extract complementary features from both modalities.
-Skeletongait++[9] employed the Skeleton Map to represent the skeleton and integrated it with the silhouette using an attention mechanism.
-Human parsing in gait recognition typically relies on representations that contain fine-grained semantic information about individual body parts.
-These methods achieve simple but efficient human body parsing based on RGB images.
-Specifically, ParsingGait[13] combined CNNs and GCNs to effectively capture both global and local gait features, thereby enhancing gait recognition performance in real-world scenarios.
-GaitParsing[14] decomposed gait into specific body parts using human semantic parsing and employed a dual-branch feature extraction network to process both holistic and part-level features.
-LandmarkGait[16] introduced an unsupervised method that converts silhouettes into consistent landmarks for fine-grained parsing, utilizing a multi-scale feature extractor to capture both global and part-level features for gait recognition.
-III Method
+2.3. Human parsing in gait recognition
+Human parsing is a specific task of semantic segmentation for pixel-level classification of human parts(Liu et al., 2014; Luo et al., 2018; Zhou et al., 2018; Wang et al., 2020; Liu et al., 2022; Gong et al., 2018, 2017).
+Human parsing in gait recognition involves segmenting a person’s body into distinct regions or parts and using these parts to enhance gait recognition accuracy.
+This fine-grained approach enables the capture of body part-specific motion patterns, which are critical for recognizing individuals under challenging conditions, such as occlusions, clothing changes, or varying walking styles.
+Human parsing techniques typically rely on semantic segmentation or human body pose estimation, which involves detecting and labeling body parts in visual representations (e.g., RGB images or silhouettes).
+Recent research in this domain has focused on leveraging both global and local features to improve gait recognition.
+ParsingGait (Zheng et al., 2023) combines convolutional neural networks (CNNs) and graph convolutional networks (GCNs) to capture both global and local gait features, thereby improving recognition performance under complex real-world conditions.
+Similarly, GaitParsing (Wang et al., 2023a) decomposes gait sequences into specific body parts using human semantic parsing.
+This approach employs a dual-branch feature extraction network to process both holistic and part-level features, allowing it to better capture fine-grained motion patterns.
+LandmarkGait (Wang et al., 2023b) introduces an unsupervised method that converts silhouettes into consistent landmarks, which can then be used for fine-grained parsing.
+This approach also incorporates a multi-scale feature extractor to capture both global and part-level features, further enhancing the gait recognition process.
+3. METHOD
 The PSGait proposed in this study is illustrated in Fig.
@@
 Finally, this multimodal gait representation, which encapsulates comprehensive information, is fed into the gait recognition model to enhance its classification performance.
-III-A Parsing Section
+3.1. Parsing Skeleton Representation
+In this section, we introduce the definition of the novel representation, i.e., Parsing Skeleton, for gait recognition.
+Given a human walking sequence of frames $S=\{\mathbf{C}^{i}\}^{N}_{i=1}$, where $\mathbf{C}^{i}\in\mathbb{R}^{K\times 2}$ is one coordinates frame and $N$ is the length of the sequence, we can feed each frame $\mathbf{C}^{i}$ into the parsing Section $P(\cdot)$ to obtain the Parsing Skeleton frame $\textbf{x}^{i}=P(\textbf{C}^{i})$.
+The parsing section $P(\cdot)$ can set $K$ body parts (including the background) with different colors for each pixel in an image of the same size as $\textbf{C}^{i}$.
+Therefore, the Parsing Skeleton sequence can be formulated as:
+$X=\{\mathbf{x}^{i}\}^{N}_{i=1},$
+where $\textbf{x}^{i}\in\mathbb{R}^{H\times W}$ is the Parsing Skeleton frame. For each pixel $p^{j}$ of $\mathbf{x}^{i}$, we have $p^{j}\in\{0,1,...,K-1\}$ where 00 denotes the background.
+Based on the above definition, we compare the proposed Parsing Skeleton representation with the binary silhouette from the viewpoint of information theory.
+Inspired by the definition of information entropy proposed by Claude E.
+Shannon (Shannon, 1948), the pixel-level information entropy of the Parsing Skeleton can be denoted as
+$\mathcal{H}=-\sum_{k=0}^{K}p_{k}\cdot log(p_{k}),$
+where pk is the possibility of one pixel belonging to the k-th class.
+For the conventional binary silhouette, each pixel only belongs to the foreground or the background, i.e., the $K=2$, so the information of the binary pixel is only one bit.
+The combination of Parsing Skeleton with silhouette, where K = 13, leads to a significantly higher information entropy compared to other gait recognition methods.
+This is due to the richer representation of 13 distinct human parts as shown in Fig.
+2, allowing for a more complete capture of gait information and enhancing the model’s ability to distinguish subtle differences in gait patterns.
+Therefore, we consider Parsing Skeleton as a gait representation with higher information entropy to encode the fine-grained dynamics of body parts during walking.
+3.2. Parsing Section
 The main objective of this section is to obtain representations with fine-grained semantic information of various body parts from the skeleton.
@@
 This results in the following mapping between human body joints $k_{i}$ and their corresponding body parts $P_{j}$:
-$h(P_{j})=\{k_{1},k_{2},\dots,k_{m}\}$ (2)
+$h(P_{j})=\{k_{1},k_{2},\dots,k_{m}\}$
 When rendering, we need to ensure that the Parsing Skeleton’s information is effectively captured by the CNN when it is fed into the gait recognition model with uniform size.
@@
 For the head representation, we introduce circles centered at the three key head points ($x,y$), with a specified radius $r$, encompassing the around the pixel $(u,v)$:
-$\text{circle}(x,y,r,c)=\{(u,v)\mid(u-x)^{2}+(v-y)^{2}\leq r^{2}\}$ (3)
+$\text{circle}(x,y,r,c)=\{(u,v)\mid(u-x)^{2}+(v-y)^{2}\leq r^{2}\}$
 For the representation of each limb, a line of specified width is parsed on the image:
-$\text{line}((x_{1},y_{1}),(x_{2},y_{2}),w,c)=\left\{(u,v)\mid d\leq\frac{w}{2}\right\}$ (4)
+$\text{line}((x_{1},y_{1}),(x_{2},y_{2}),w,c)=\left\{(u,v)\mid d\leq\frac{w}{2}\right\}$
 where $d$ represents the distance of the pixel $(u,v)$ from the line, $c$ denotes the color assigned to this part, and $w$ is the width of this line.
-III-B Fusion Strategy
+3.3. Fusion Strategy
 In this study, we investigate two distinct fusion strategies for combining Parsing Skeleton and silhouette in representation learning.
@@
 Specifically, the fused image $F_{\text{overlay}}(x,y)$ is defined as follows:
-$F_{\text{overlay}}(x,y)=\max\left(M(x,y),S(x,y)\right),$ (5)
+$F_{\text{overlay}}(x,y)=\max\left(M(x,y),S(x,y)\right)$
 where $M(x,y)$ is the binary silhouette image, and $S(x,y)$ represents the Parsing Skeleton.
@@
 The overlay method produces an RGB fused image, where the silhouette and Parsing Skeleton share the same spatial space, allowing the network to jointly learn the global shape features and structural details with minimal processing complexity.
+Figure 3.
+Visualization of different gait representations.
+(a) Raw RGB frames from a walking sequence.
+(b) Binary silhouette images extracted via background subtraction.
+(c) 2D skeleton keypoints estimated from the RGB frames.
+(d) Our proposed Parsing Skeleton, which integrates part-level semantics and structural connectivity, providing richer spatial and motion cues for gait recognition.
+(e) Parsing Skeleton combined with silhouettes — the final fused representation used in PSGait, combining part-level semantics and body contours to enhance discriminability and structural consistency in gait recognition.
 Multi-Channel Fusion: In the second strategy, we introduce a more granular representation by separating each Parsing Skeleton component (lines and circles) into independent channels.
@@
 This results in a multi-channel representation $F_{\text{multi}}$, where the fused feature map has 12 channels:
-$F_{\text{multi}}(x,y,c)=\begin{cases}S_{c}(x,y),&\text{if }c=1,2,\dots,11\\ M(x,y),&\text{if }c=12\end{cases}$ (6)
+$F_{\text{multi}}(x,y,c)=\begin{cases}S_{c}(x,y),&\text{if }c=1,2,\dots,11\\ M(x,y),&\text{if }c=12\end{cases}$
 where $S_{c}(x,y)$ represents the $c$-th Parsing Skeleton component (either a line or a circle), and the 12th channel contains the silhouette image.
 While this strategy provides a more explicit and structured representation, enabling the network to focus on individual components of the Parsing Skeleton and silhouette, it inherently increases the complexity of the input and may dilute spatial relationships across channels.
-III-C Align Center
-After the data fusion stage, the center alignment operation is performed to ensure that the human body contour and the resolved skeleton are centered in the image under the premise of uniform input image size.
-This alignment facilitates subsequent model processing and feature extraction.
-Given a binary silhouette image $M(x,y)$, where $M(x,y)\in\{0,1\}$, the upper bound $y_{\text{top}}$ and lower bound $y_{\text{btm}}$ of the human body region are calculated as follows:
-$y_{\text{top}}=\min\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (7)
-$y_{\text{btm}}=\max\left\{y\,\middle|\,\sum_{x}M(x,y)>0\right\}.$ (8)
-Next, the image is cropped based on the upper and lower bounds of the human region and resized to the target height while preserving the aspect ratio, resulting in the image $M^{\prime}(x,y)$.
-The center of the horizontal axis is calculated as follows:
-$x_{c}=\frac{\sum_{x,y}x\cdot M^{\prime}(x,y)}{\sum_{x,y}M^{\prime}(x,y)}$ (9)
-Using the horizontal axis center $x_{c}$, the image is cropped to a target size of $H^{\prime}\times W^{\prime}$.
-TABLE I: Quantitative comparison of gait recognition methods on the SUSTech1K[19] is presented.
+3.4. Data Augmentation
+To improve generalization and reduce overfitting, we apply data augmentation during training.
+Since gait sequences are already spatially aligned, we adopt a general image augmentation strategy following (Fan et al., 2023b), including: (1)Horizontal Flipping.
+Simulates varying walking directions without altering gait features, enhancing robustness to orientation changes.
+(2)Affine and Perspective Transformations.
+Introduces minor viewpoint variations through random geometric distortions, improving model adaptability to different camera angles.
+(3)Random Erasure.
+Randomly occludes image regions to mimic occlusion and background noise, encouraging the model to rely on global rather than local features.
+These strategies diversify the training data and improve the model’s robustness across real-world conditions.
+Table 1.
+Quantitative comparison of gait recognition methods conducted across two authoritative datasets, CCPG(Li et al., 2023b) and Gait3D(Zheng et al., 2022).
 The best performances are indicated in bold, and the second-best methods are underlined.
 †We modified Skeletongait++ to match the settings used in GaitBase.
-Input Method Probe Sequence(R-1) Overall
-Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
-Skeleton GaitGraph2(CVPRW2022)[20] 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
-GaitTR(ES2023)[21] 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
-Silhouette GaitSet(AAAI2019)[22] 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
-GaitPart(CVPR2019)[23] 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
-GaitGL(ICCV2021)[24] 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
-GaitBase(CVPR2023)[18] 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
-DeepGaitV2(Arxiv2023)[25] 83.5 79.5 46.3 76.8 79.1 78.5 81.1 27.3 77.4 90.2
-[…]
-Skeleton+Silhouette BiFusion(MTA2023)[10] 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
-SkeletonGait++†(AAAI2024)[9] 84.7 82.4 46.7 81.8 81.2 82.6 85.1 44.3 81.2 93.2
-[…]
-[…]
-TABLE II: Quantitative comparison of gait recognition methods is conducted across two authoritative datasets, CCPG[17] and Gait3D[8].
-†We modified Skeletongait++ to match the settings used in GaitBase.
-Input Method Testing Datasets
+Input Method Source Testing Datasets
 CCPG Gait3D
 CL UP DN BG Mean rank-1 rank-5 mAP mINP
-Skeleton GaitGraph2(CVPRW2022)[20] 5.0 5.3 5.8 6.2 5.6 11.1 - - -
-GaitTR(ES2023)[21] 15.7 18.3 18.5 17.5 17.5 6.6 - - -
-Silhouette GaitSet(AAAI2019)[22] 60.2 65.2 65.1 68.5 64.8 36.7 58.3 30.0 17.3
-GaitPart(CVPR2020)[23] 64.3 67.8 68.6 71.7 68.1 28.2 47.6 21.6 12.4
-GaitBase(CVPR2023)[18] 71.6 75.0 76.8 78.6 75.5 64.6 - - -
-[…]
-[…]
-Skeleton+Silhouette SkeletonGait++†(AAAI2024)[9] 77.0 81.4 83.0 88.8 82.6 64.8 80.5 54.8 36.1
-[…]
-[…]
-IV Experience
-IV-A Datasets
-We evaluated our proposed method on three commonly used datasets: SUSTech1K[19] and CCPG[17], which are collected in fully constrained laboratory environments, and Gait3D[8], which is collected from real-world scenarios.
-SUSTech1K[19] contains gait data from 1,050 subjects, covering 8 different conditions: normal (NM), bag (BG), clothes changing (CL), carrying (CR), umbrella (UB), uniform (UF), occlusion (OC), and night (NT), as well as 12 viewpoints (ranging from 0° to 315°).
-The diversity and granularity of the dataset make SUSTech1K[19] an ideal platform for studying the performance of gait recognition in complex environments.
-CCPG[17] was designed to address the challenge of gait recognition under clothing change conditions and is particularly suitable for dynamic person recognition in video surveillance.
+Skeleton GaitGraph2(Teepe et al., 2022) CVPRW2022 5.0 5.3 5.8 6.2 5.6 11.1 -
+GaitTR(Zhang et al., 2023) ES2023 15.7 18.3 18.5 17.5 17.5 6.6 -
+Silhouette GaitSet(Chao et al., 2021) AAAI2019 60.2 65.2 65.1 68.5 64.8 36.7 58.3 30.0 17.3
+GaitPart(Fan et al., 2020) CVPR2020 64.3 67.8 68.6 71.7 68.1 28.2 47.6 21.6 12.4
+GaitBase(Fan et al., 2023b) CVPR2023 71.6 75.0 76.8 78.6 75.5 64.6 -
+DeepGaitV2(Fan et al., 2023a) Arxiv2023 78.6 84.8 80.7 89.2 83.3 74.4 88.0 65.8 -
+Silhouette+ Skeleton SkeletonGait++(Fan et al., 2024) AAAI2024 77.0 81.4 83.0 88.8 82.6 64.8 80.5 54.8 36.1
+PSGait-B Ours 82.5 85.3 88.6 89.1 86.4 70.0 85.7 62.2 43.3
+PSGait-D Ours 80.9 84.8 86.4 90.1 85.6 76.2 88.7 69.4 50.8
+Table 2. Quantitative comparison of gait recognition methods on the SUSTech1K(Shen et al., 2023).
+Input Method Publication Probe Sequence (R-1) Overall
+Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
+Skeleton GaitGraph2(Teepe et al., 2022) CVPRW2022 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
+GaitTR(Zhang et al., 2023) ES2023 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
+Silhouette GaitSet(Chao et al., 2021) AAAI2019 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
+GaitPart(Fan et al., 2020) CVPR2019 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
+GaitGL(Lin et al., 2022) ICCV2021 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
+GaitBase(Fan et al., 2023b) CVPR2023 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
+DeepGaitV2(Fan et al., 2023a) Arxiv2023 83.5 79.5 46.3 76.8 79.1 78.5 81.1 27.3 77.4 90.2
+Silhouette+ Skeleton BiFusion(Peng et al., 2024) MTA2023 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
+SkeletonGait++(Fan et al., 2024) AAAI2024 84.7 82.4 46.7 81.8 81.2 82.6 85.1 44.3 81.2 93.2
+PSGait-B Ours 86.2 84.5 55.1 83.0 85.4 84.2 90.7 47.2 83.4 94.3
+PSGait-D Ours 87.9 85.2 53.9 84.5 86.3 85.3 91.4 47.5 84.6 94.8
+3.5. Gait Recognition Model
+The current gait recognition framework employs a backbone network $\mathcal{B}$ to extract features $\mathbf{f}$ from an input sequence of gait images $\mathbf{s}$, defined as:
+$\mathbf{f}=\mathcal{B}(\mathbf{s})\in\mathbb{R}^{n\times c\times s\times h\times w}$
+where $n$, $c$, $s$, $h$, and $w$ denote the number of gait frames, the number of channels, the sequence length, and the height and width of each image, respectively.
+Temporal Pooling (TP) (Chao et al., 2021) is applied to compress these features into more compact representations $z$:
+$z=TP({\mathbf{f}})\in\mathbb{R}^{n\times c\times h\times w}$
+Subsequently, Horizontal Pooling (HP) (Fu et al., 2019) divides these feature maps into horizontal segments, pooling each segment into vectors $z_{s}$ to enable efficient feature extraction through global pooling across these segments:
+$\mathbf{f}^{\prime}=\text{maxpool}(z_{s})+\text{avgpool}(z_{s})$
+A fully connected layer then projects these pooled vectors into a metric space, followed by Batch Normalization Neck (BNNeck) (Luo et al., 2019) to refine the feature space.
+The final classification is performed via cross-entropy loss:
+$L_{ce}=-\sum_{i=1}^{n}y_{i}\log(\widehat{y_{i}})$
+where $y_{i}$ denotes the true label and $\widehat{y_{i}}$ represents the predicted probability for each class.
+Additionally, the model employs triplet loss to enhance the discrimination of subtle gait variations essential for accurate classification.
+During training, $N$ triplets are formed in each batch, where each triplet consists of an anchor sequence $\mathbf{s}_{i}^{a}$, a positive sequence $\mathbf{s}_{i}^{p}$ from the same identity, and a negative sequence $\mathbf{s}_{i}^{n}$ from a different identity.
+The triplet loss(Schroff et al., 2015) is computed as follows:
+$L_{triplet}=\sum_{i=1}^{N}\max\left(0,\|f(\mathbf{s}_{i}^{a})-f(\mathbf{s}_{i}^{p})\|_{2}^{2}-\|f(\mathbf{s}_{i}^{a})-f(\mathbf{s}_{i}^{n})\|_{2}^{2}+\alpha\right)$
+where $f(\cdot)$ is the embedding function that maps each sequence to the embedding space, $|\cdot|_{2}$ represents the Euclidean distance, and $\alpha$ is a margin parameter that enforces a separation between positive and negative pairs.
+This strategy improves the model’s discriminative ability between classes.
+The overall objective function is a combination of the cross-entropy loss and the triplet loss, formulated as:
+$L_{total}=L_{ce}+L_{triplet}$
+This integrated loss function optimizes the gait recognition model, enhancing its accuracy in distinguishing between different identities based on gait features.
+Table 3. Generalization experiments on two baseline models across three datasets.
+Method SUSTech1K Gait3D CCPG
+rank-1 rank-5 rank-1 mAP rank-1
+GaitBase(Fan et al., 2023b) 76.1 89.4 64.6 54.5 75.5
+GaitBase+Parsing Skeleton 83.4(7.3 $\uparrow$) 94.3(4.9 $\uparrow$) 70.0(5.4 $\uparrow$) 62.2(7.7 $\uparrow$) 86.4(10.9$\uparrow$)
+DeepGaitV2(Fan et al., 2023a) 77.4 90.2 74.4 65.8 83.3
+DeepGaitV2+Parsing Skeleton 84.6(7.2 $\uparrow$) 94.8(4.6 $\uparrow$) 76.2(1.8 $\uparrow$) 69.4(3.6$\uparrow$) 85.6(2.3$\uparrow$)
+Table 4. Efficiency and accuracy comparison between PSGait-B and the SOTA method.
+Method Parameters GPU memory Preprocessing time Training time SUSTech1K CCPG
+(M) (GB) (ms per frame) (s per epoch) rank-1 rank-1
+GaitBase(Fan et al., 2023b) 8.02 4.68 0 0.22 76.1 75.5
+SkeletonGait++(Fan et al., 2024) 9.13 5.14 0.71 0.34 81.3 83.7
+PSGait-B(Ours) 8.02(12$\downarrow$) 4.70(9$\downarrow$) 0.34(52$\downarrow$) 0.27(21$\downarrow$) 83.4(2$\uparrow$) 86.4(3$\uparrow$)
+4. EXPERIMENTS
+4.1. Datasets
+We evaluated our proposed method on three commonly used datasets: SUSTech1K(Shen et al., 2023) and CCPG(Li et al., 2023b), which are collected in fully constrained laboratory environments, and Gait3D(Zheng et al., 2022), which is collected from real-world scenarios.
+Our experiments strictly follow the official evaluation protocols.
+SUSTech1K(Shen et al., 2023) contains gait data from 1,050 subjects, covering 8 different conditions: normal (NM), bag (BG), clothes changing (CL), carrying (CR), umbrella (UB), uniform (UF), occlusion (OC), and night (NT), as well as 12 viewpoints (ranging from 0° to 315°).
+The diversity and granularity of the dataset make SUSTech1K(Shen et al., 2023) an ideal platform for studying the performance of gait recognition in complex environments.
+CCPG(Li et al., 2023b) was designed to address the challenge of gait recognition under clothing change conditions and is particularly suitable for dynamic person recognition in video surveillance.
 The dataset includes 200 identities and more than 16,000 sequences, with each identity captured from 10 viewpoints under seven different clothing change conditions, covering both indoor and outdoor environments.
 It provides two data modalities: RGB images and silhouette images.
-Gait3D[8] is a large-scale gait dataset collected in real-world environments.
+Gait3D(Zheng et al., 2022) is a large-scale gait dataset collected in real-world environments.
 It was recorded in a supermarket and contains 1,090 hours of video footage with a resolution of 1,920×1,080 and a frame rate of 25 FPS.
 Gait3D provides four modalities: silhouettes, 2D and 3D skeletons, and 3D meshes.
-TABLE III: Impact of circle radius and line width on the SUSTech1K[19] dataset using GaitBase[18].
+4.2. Implementation Details
+For data preprocessing, silhouettes are extracted from all three datasets.
+Since skeleton data is unavailable for CCPG, we employ Sapiens (Khirodkar et al., 2024) for 2D pose estimation, while original skeleton annotations are used for SUSTech1K and Gait3D in COCO17(Lin et al., 2014) format, with a confidence threshold of 0.1.
+In the parsing stage, to ensure the Parsing Skeleton is clearly captured by CNNs, the circle radius (for the head) and line width (for limbs) are adjusted according to image size.
+Specifically, SUSTech1K uses radius 10 and width 12; CCPG and Gait3D use radius 16 and width 20.
+The color scheme for each body part is shown in Fig.
+2.
+To evaluate generalization, Parsing Skeletons are integrated into two state-of-the-art gait recognition models: GaitBase(Fan et al., 2023b) and DeepGaitV2(Fan et al., 2023a).
+The fused inputs are resized to $64\times 44$ before being fed into the recognition models.
+For training, we almost follow the OpenGait (Fan et al., 2023b) framework.
+On SUSTech1K(Shen et al., 2023), batch size is set to (8, 8), with learning rate milestones at 20k, 30k, and 40k steps, and a total of 50k steps.
+On CCPG(Li et al., 2023b), we use (8, 16) with milestones at 20k, 40k, and 50k (60k total).
+On Gait3D(Zheng et al., 2022), the batch size is (32, 4), also trained for 60k steps.
+We use SGD with an initial learning rate of 0.1 and momentum of 0.9 to facilitate stable convergence.
+Table 5. Impact of circle radius and line width on the SUSTech1K(Shen et al., 2023) dataset using GaitBase(Fan et al., 2023b).
 Circle Radius Line Width Probe Sequence (R-1) Overall
@@
 3 3 82.1 80.2 47.0 77.9 80.8 79.4 87.1 33.8 78.8 91.7
-10 12 86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
+10 12 86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 47.1
 20 24 84.6 83.1 51.3 80.8 83.0 81.2 90.0 43.1 81.6 93.7
-TABLE IV: Impact of input modalities and fusion strategies on the SUSTech1K[19] dataset using GaitBase[18].
-Skeleton Silehouette Overlay Probe Sequence(R-1) Overall
-Normal Bag Clothing Carrying Umbrella Uniform Occlusion Night R-1 R-5
-52.5 50.1 35.2 48.4 40.3 51.3 61.1 22.0 50.3 72.5
-81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
-82.7 81.2 47.8 78.9 81.3 79.8 88.2 35.1 79.7 92.2
-86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
-IV-B Implementation Details
-For data preprocessing, the original data is used to extract silhouette information from all three datasets.
-Since skeleton data is not available for the CCPG dataset, we utilize the Sapiens[26] method for pose estimation.
-The original skeleton data is used for the other two datasets, and the format of points follows COCO17.
-A confidence threshold of 0.1 is applied.
-In the parsing stage, to ensure that the Parsing Skeleton features of the processed image can be effectively captured by the CNN, the radius of the circle representing the head and the width of the line segment representing the limbs are determined based on the image size ratio in each dataset.
-Specifically, for the SUSTech1K[19] dataset, a radius of 10 and a line width of 12 are used.
-For the CCPG[17] dataset, the radius and line width are set to 16 and 20, respectively.
-Similarly, for the Gait3D[8] dataset, the radius and line width are also set to 16 and 20, respectively.
-The colors selected for parsing each body part are shown in Fig.
-2.
-To evaluate the versatility of our method, we apply it to two state-of-the-art gait recognition models: GaitBase[18] and DeepGaitV2[25].
-The size of the fused image is adjusted to $64\times 44$ before being input into the gait recognition model.
-For model training, we largely follow the parameter settings from OpenGait[18], an excellent open-source framework in the field of gait recognition.
-Specifically, for the SUSTech1K[19] dataset, the batch size is set to $(8,8)$, and learning rate milestones are scheduled at 20k, 30k, and 40k steps, with a total of 50k training steps.
-For the CCPG[17] dataset, we use a batch size of $(8,16)$ with learning rate milestones at 20k, 40k, and 50k steps, and a total of 60k iterations.
-For the Gait3D[8] dataset, the batch size is set to $(32,4)$, with learning rate milestones occurring at 20k, 40k, and 50k steps, and a total of 60k iterations.
-The optimization process is configured using the SGD solver.
-Specifically, the learning rate $lr$ is initialized at 0.1, providing an appropriate step size for updating the model weights during training.
-A momentum value of 0.9 is employed to accelerate convergence by reducing oscillations and incorporating a weighted history of previous gradients.
-IV-C Comparison with State-of-the-art Methods
-As shown in Table I and Table II, we evaluate the performance on three datasets. The results highlight the strong competitiveness and versatility of our method.
-Competitiveness.
-We compare and validate the advantages of our method across three datasets.
-On the SUSTech1K[19] dataset (Table I), compared to SkeletonGait++, Ours+GaitBase improves the Rank-1 accuracy from 81.2% to 83.4% (+2.2%).
-This demonstrates the effectiveness of our method under the same baseline settings.
-Furthermore, Ours+DeepGaitV2 achieves the best overall performance on SUSTech1K, attaining a Rank-1 accuracy of 87.9% and a Rank-5 accuracy of 94.8%.
-On the CCPG[17] dataset (Table II), compared to SkeletonGait++[9], Ours+GaitBase improves the mean accuracy from 82.6% to 86.4% (+3.8%).
-On the Gait3D[8] dataset (Table II), Rank-1 accuracy is improved from 64.8% to 70.0% (+5.2%).
-Additionally, Ours+GaitBase and Ours+DeepGaitV2 achieve the best overall performance on CCPG and Gait3D, with a mean accuracy of 86.4% on CCPG and a Rank-1 accuracy of 85.6% and mAP of 69.4% on Gait3D, outperforming all other state-of-the-art methods.
-These improvements underscore the robustness and adaptability of our approach across diverse benchmarks and evaluation metrics.
-Versatility.
-We further demonstrate the versatility of our approach by integrating it into two baseline models, GaitBase[18] and DeepGaitV2[25], and comparing the resulting performance improvements.
-On the three datasets mentioned above, our method provides a significant boost over the existing baselines.
-Specifically, on the SUSTech1K dataset[19], Ours+GaitBase improves the Rank-1 accuracy from 76.1% to 83.4% (+7.3%), while Ours+DeepGaitV2 enhances the Rank-1 accuracy from 77.4% to 84.6% (+7.2%).
-On the CCPG[17] dataset, integrating our method with GaitBase increases the mean accuracy from 75.5% to 86.4% (+10.9%).
-For DeepGaitV2, the mean accuracy improves from 83.3% to 85.6% (+2.3%).
-On Gait3D[8], Ours+GaitBase improves the Rank-1 accuracy from 64.6% to 70.0% (+5.4%).
-For DeepGaitV2, the Rank-1 accuracy increases from 74.4% to 76.2% (+1.8%), and the mAP rises from 65.8% to 69.4% (+3.6%).
-These consistent performance improvements across multiple datasets and baseline models highlight the strong versatility of our method.
-It effectively enhances the performance of any gait recognition model, making it plug-and-play and universally applicable to various scenarios and datasets.
-IV-D Ablation Study
-To validate the effectiveness of our method, we conducted ablation studies on the SUSTech1K[19] dataset using GaitBase[18] as the baseline gait recognition model.
+Table 6. Impact of input modalities and fusion strategies on the CCPG(Li et al., 2023b) dataset using GaitBase(Fan et al., 2023b).
+Skeleton Silehouette Multi-Channel Overlay-Based CCPG
+CL UP DN BG Mean
+53.8 58.5 61.5 56.0 57.4
+76.5 80.5 82.3 86.4 81.4
+79.3 83.3 85.2 86.5 83.6
+82.5 85.3 88.6 89.1 86.4
+4.3. Comparison with State-of-the-art Methods
+We evaluate the proposed PSGait framework against state-of-the-art gait recognition methods across three input modalities-skeletons, silhouettes, and their fusion—on three benchmarks: CCPG(Li et al., 2023b), SUSTech1K(Shen et al., 2023), and Gait3D(Zheng et al., 2022).
+PSGait-B and PSGait-D adopt GaitBase (Fan et al., 2023b) and DeepGaitV2 (Fan et al., 2023a) as gait recognition models, respectively.
+In order to highlight the advantages of the Parsing Skeleton representation itself, rather than the influence of the network architecture, we adjusted the downstream settings of SkeletonGait++ to align with those of GaitBase.
+On CCPG and Gait3D (Table 1), PSGait consistently outperforms existing methods in both unimodal and multimodal settings.
+Under the skeleton-only modality, methods like GaitGraph2 (Teepe et al., 2022) and GaitTR (Zhang et al., 2023) achieve low recognition rates due to the sparsity of keypoints.
+SkeletonGait++ (Fan et al., 2024) improves performance by using Skeleton Maps, but still struggles in real-world scenarios.
+In contrast, when our Parsing Skeleton is fused with silhouettes, PSGait-B and PSGait-D set new state-of-the-art results, achieving up to 86.4% Rank-1 on CCPG and 76.2% on Gait3D.
+PSGait-D also brings large gains in mAP and mINP (+12.7% and +14.7%) over the strongest baseline, indicating improved retrieval robustness.
+On SUSTech1K (Table 2), PSGait again demonstrates superior performance under various challenging conditions.
+When our Parsing Skeleton is fused with silhouettes, PSGait-B and PSGait-D consistently outperform SkeletonGait$+$ across all probe types.
+For example, PSGait-D achieves 91.4% Rank-1 under occlusion and 86.1% overall Rank-1 accuracy, surpassing the baseline by large margins—+7.6% in night and +7.1% in carrying conditions.
+These results validate that: (1) Parsing Skeleton provides a richer and more discriminative representation than traditional silhouettes or skeletons; (2) PSGait effectively captures both part-level dynamics and global structure, enabling consistent SOTA performance across datasets; (3) PSGait is accurate, lightweight, and robust, making it highly suitable for real-world gait recognition applications.
+4.4. Generalization of the Parsing Skeleton
+To evaluate the generalization capability of the proposed Parsing Skeleton, we integrate it into two representative baselines—GaitBase and DeepGaitV2—and test on three diverse datasets: SUSTech1K, Gait3D, and CCPG (Table 3).
+On SUSTech1K, Parsing Skeleton brings notable Rank-1 improvements of +7.3% for GaitBase and +7.2% for DeepGaitV2, with consistent gains in Rank-5, demonstrating enhanced retrieval performance.
+On the challenging Gait3D dataset, GaitBase+Parsing Skeleton improves by +5.4% in Rank-1 and +7.7% in mAP, while DeepGaitV2 gains +1.8% and +3.6%, respectively—showing its effectiveness under real-world, multi-view conditions.
+On CCPG, Parsing Skeleton improves GaitBase by a substantial +10.9% and DeepGaitV2 by +2.3% in Rank-1, highlighting its robustness across varied environments.
+These results confirm that Parsing Skeleton is a model-agnostic, plug-and-play, generalizable representation.
+It consistently improves performance across architectures and datasets, making it well-suited for real-world deployment in gait recognition systems.
+4.5. Lightweight of the PSGait
+To evaluate computational efficiency, we compare PSGait-B with two baselines—GaitBase(Fan et al., 2023b) (silhouette-only) and SkeletonGait++(Fan et al., 2024) (skeleton map + silhouette)—in terms of model size, memory usage, preprocessing/training time, and accuracy.
+All evaluations were conducted on 100,000 frames from SUSTech1K(Shen et al., 2023), using GaitBase as the backbone and four RTX 2080 GPUs.
+As shown in Table 4, PSGait-B achieves a superior trade-off between efficiency and performance.
+It reduces parameters by 12%, GPU memory by 9%, preprocessing time by 52%, and training time by 21% compared to SkeletonGait++.
+Despite these reductions, PSGait-B achieves +2% Rank-1 accuracy on SUSTech1K and +3% on CCPG(Li et al., 2023b), showing better performance at lower cost.
+Compared to GaitBase, PSGait-B maintains similar resource usage but brings substantial accuracy gains (+7.3% on SUSTech1K, +10.9% on CCPG), validating that the Parsing Skeleton effectively enhances model capacity without increasing complexity.
+In summary, PSGait is a lightweight and efficient solution that delivers high performance and is well-suited for real-time applications or limited-resource environments.
+4.6. Ablation Study
+To validate the effectiveness of our method, we conducted ablation studies on SUSTech1K(Shen et al., 2023) and CCPG(Li et al., 2023b) using GaitBase(Fan et al., 2023b) as the baseline gait recognition model.
 We conducted experiments to assess the impact of key parameters, including circle radius, line width, input modalities, and different data fusion strategies.
-Influence of Circle Radius and Line Width.
-As shown in Table III, we evaluated the impact of circle radius and line width on recognition performance.
-A smaller setting results in a relatively low Rank-1 accuracy of 82.1%, while a moderate setting achieves the best Rank-1 accuracy of 86.1% and Rank-5 accuracy of 94.3%.
-However, further increasing the radius and width, slightly reduces Rank-1 accuracy to 84.6%, indicating that finding an optimal balance is crucial for performance.
-The results indicate that choosing an appropriate circle radius and line width is critical for effectively extracting parsing skeleton features.
-If these values are too small, the parsing skeleton cannot be effectively captured by the CNN after resizing the image.
-Conversely, if they are too large, the parsing skeleton may obscure the silhouette, impeding the extraction of human shape features by gait recognition models.
-Impact of Data Fusion Strategies.
-We further analyzed the contribution of different modalities and fusion methods (parsing skeleton, silhouette, and overlay) to overall performance.
-As shown in Table IV, using parsing skeletons or silhouettes alone does not yield particularly high performance, indicating that relying on either type of data alone is insufficient for accurate gait recognition.
-Incorporating silhouette information significantly improves performance, with Rank-1 increasing to 79.7% and Rank-5 reaching 92.2%.
-In this case, the data fusion strategy employed is a multi-channel fusion method.
-When parsing skeleton and silhouette are overlaid, the model achieves the best performance, with a Rank-1 accuracy of 83.4% and a Rank-5 accuracy of 94.3%.
-The overlay fusion strategy capitalizes on the complementary strengths of skeleton and silhouette data, allowing the model to capture both global shape and local structural details.
-These results demonstrate that our fusion strategy is both effective and achieves an optimal balance between accuracy and robustness.
-V Conclusion
-In this paper, we propose a novel gait representation, Parsing Skeleton, which leverages skeleton-guided human parsing to capture fine-grained body dynamics.
+Circle Radius and Line Width.
+As shown in Table 5, we evaluated the impact of circle radius and line width on recognition performance.
+A smaller setting leads to relatively lower accuracy, while a moderate setting yields the best performance in terms of Rank-1 accuracy.
+However, increasing the radius and width beyond this point results in a slight decrease in Rank-1 accuracy.
+Smaller values lead to under-represented structures, while larger ones may obscure silhouette contours, degrading performance.
+This highlights the importance of carefully balancing visual clarity and structural integrity when rendering Parsing Skeletons.
+Fusion Strategies.
+Table 6 compares different input modalities and fusion approaches.
+Here, “Skeleton” refers to Parsing Skeleton.
+Using Parsing Skeletons or silhouettes alone does not yield particularly high performance, indicating that relying on either type of data alone is insufficient for accurate gait recognition.
+Incorporating skeleton and silhouette information improves performance, with Rank-1 increasing to 83.6%.
+The best performance is achieved by overlaying Parsing Skeletons onto silhouettes.
+This overlay strategy effectively integrates part-level dynamics with global shape cues, enabling more robust and discriminative gait representations.
+These results confirm that both visual rendering parameters and fusion strategy are crucial to fully unlocking the potential of Parsing Skeleton for gait recognition.
+5. CONCLUSION
+In this paper, we propose a novel gait representation, Parsing Skeleton, which not only captures fine-grained body dynamics through skeleton-guided human parsing, but also transforms skeleton data into image form, enabling effective feature extraction via CNNs.
 This representation addresses the limitations of traditional gait recognition methods that rely on silhouettes or skeletons with low information entropy, by encoding more detailed and comprehensive features of human motion.
 To effectively exploit the Parsing Skeleton, we propose PSGait, a new gait recognition framework that integrates Parsing Skeletons with silhouettes to enhance feature robustness and improve individual differentiation, particularly in unconstrained, real-world scenarios.
-Extensive experiments across various datasets demonstrate that PSGait outperforms existing state-of-the-art multimodal gait recognition methods.
+Extensive experiments across various datasets demonstrate that PSGait outperforms existing state-of-the-art multimodal gait recognition methods while significantly reducing computational resources.
 As a plug-and-play method, PSGait leads to significant performance improvements across various models.
-This work validates the effectiveness and versatility of the parsing skeleton representation for real-world gait recognition tasks, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
+This work validates the effectiveness, generalization and lightweight of the Parsing Skeleton representation for real-world gait recognition tasks, establishing PSGait as a new state-of-the-art approach for multimodal gait recognition.
+We believe that Parsing Skeleton, as a general human representation, may also provide inspiration for other domains involving human motion analysis.
+Appendix A Supplementary Material
+In this section, we provide the following supplementary materials that are not included in our main paper due to space constraints:
+  • Parsing Skeleton generation details.
+  • Detailed dataset configuration.
+  • Additional experimental details.
+  • Gait recognition model architecture details.
+  • Limitation.
+  • Future work.
+A.1. Parsing Skeleton Generation Details
+To extract a more semantically meaningful representation of human gait, we introduce a Parsing Skeleton formulation, as illustrated in Figure 4.
+The process begins with pose estimation outputs in the form of joint coordinates (a).
+These keypoints are then converted into image-based skeleton representation by connecting them with geometric primitives—circles at joints and lines for limbs (b).
+The purpose of this step is to ensure that the image-based skeleton representation can be directly captured by CNNs.
+To further enrich the representation, we assign distinct semantic colors to each anatomical part, including the ears, arms, thighs, and calves, producing what we refer to as the Parsing Skeleton (c).
+This crucial step incorporates part-level identity, allowing for a more refined and detailed structural understanding.
+Finally, the Parsing Skeleton is fused with the silhouette image (d), resulting in a hybrid representation that combines both pose dynamics and body shape information.
+This fusion not only enhances spatial completeness but also provides a stronger foundation for downstream gait recognition tasks.
+Table 7. Datasets in use. #ID and #Seq present the number of identities and sequences.
+DataSet Train Set Test Set
+Id Seq Id Seq
+CCPG 100 8,187 100 8,095
+SUSTech1K 200 5,988 850 19,228
+Gait3D 3,000 18,940 1,000 6,369
+A.2. Detailed Dataset Configuration
+Table 7 summarizes the statistics of the datasets used in our experiments, including the number of identities and sequences in both training and testing sets.
+Our experiments strictly follow the official evaluation protocols.
+Specifically, the CCPG(Li et al., 2023b) dataset contains 100 identities in both training and testing sets, with 8,187 and 8,095 sequences, respectively.
+The SUSTech1K(Shen et al., 2023) dataset includes 200 identities and 5,988 sequences for training, and 850 identities with 19,228 sequences for testing.
+Gait3D(Zheng et al., 2022) is the largest among the three, comprising 3,000 identities and 18,940 sequences in the training set, and 1,000 identities with 6,369 sequences in the testing set.
+These datasets provide a diverse and comprehensive benchmark for evaluating gait recognition performance.
+Figure 4.
+Illustration of the Parsing Skeleton construction process.
+(a) Original pose keypoints in coordinate form.
+(b) Skeleton image generated by connecting joints with lines and circles.
+(c) Parsing skeleton with semantic part colors.
+(d) Final Parsing Skeleton fused with silhouette, enriching body structure details with spatial and semantic cues.
+A.3. Additional Experimental Details
+Table 8 presents the implementation details for training on the three datasets.
+The batch size is denoted as a tuple $(q,k)$, where $q$ is the number of subjects and $k$ is the number of sequences sampled per subject.
+For CCPG(Li et al., 2023b), the batch size is (8, 16), and the learning rate is decayed at 20k, 40k, and 50k steps, with training ending at 60k steps.
+SUSTech1K(Shen et al., 2023) uses a batch size of (8, 8), with milestones at 20k, 30k, and 40k, and a total of 50k training steps.
+Gait3D(Zheng et al., 2022) adopts a batch size of (32, 4), with learning rate milestones identical to those of CCPG, and is trained for 60k steps in total.
+Table 8. Implementation details. The batch size $(q,k)$ indicates $q$ subjects with $k$ sequences per subject.
+DataSet Batch Size Milestones Total Steps
+CCPG (8, 16) (20k, 40K, 50k) 60k
+SUSTech1K (8, 8) (20k, 30k, 40k) 50k
+Gait3D (32, 4) (20k, 40K, 50k) 60k
+Table 9. Architectures of DeepGaitV2. Down-sampling is performed by Stage2 and Stage3 with a stride of 2. $T$ and $C$ respectively denote the input sequence length and arbitrary channel number.
+Layers
+Output feature
+map size
+Block structure
+Conv0 ($T$, $C$, 64, 44) $3\times 3$, stride 1
+Stage1 ($T$, $C$, 64, 44) $\begin{bmatrix}3\times 3,C\\ 3\times 3,C\end{bmatrix}$
+Stage2 ($T$, $2C$, 32, 22) $\begin{bmatrix}3\times 3\times 3,2C\\ 3\times 3\times 3,2C\end{bmatrix}$
+Stage3 ($T$, $4C$, 16, 11) $\begin{bmatrix}3\times 3\times 3,4C\\ 3\times 3\times 3,4C\end{bmatrix}$
+Stage4 ($T$, $8C$, 16, 11) $\begin{bmatrix}3\times 3\times 3,8C\\ 3\times 3\times 3,8C\end{bmatrix}$
+TP (1, $8C$, 16, 11) Temporal Pooling
+HP (1, $8C$, 16, 1) Horizontal Pooling
+Head
+(1, $8C$, 16, 1) Flatten, 16 separate fully-connected layers and BNNecks
+A.4. Gait Recognition Model Architecture Details
+PSGait-D uses DeepGaitV2(Fan et al., 2023a) as the gait recognition model.
+Table 9 summarizes the architecture of the DeepGaitV2-3D .
+The model takes input sequences of size $(T,C,64,44)$ and progressively extracts features through four stages, with down-sampling applied at Stage2 and Stage3.
+Spatiotemporal convolutions ($3\times 3\times 3$) are used in deeper layers to capture temporal dynamics.
+Temporal and horizontal pooling are then applied, followed by a head with 16 parallel fully-connected layers and BNNecks for final representation learning.
+PSGait-B adopts GaitBase(Fan et al., 2023b) as the gait recognition model, which is built upon a ResNet-9 backbone.
+A.5. Limitation
+While PSGait demonstrates strong performance in gait recognition tasks, it still encounters several challenges under complex real-world conditions.
+Specifically, its recognition accuracy tends to degrade in the presence of extreme occlusion, drastic clothing variations, and unfavorable lighting conditions.
+Although the Parsing Skeleton representation captures fine-grained body dynamics and provides more detailed structural information, it remains limited when critical body parts are heavily occluded or entirely invisible.
+Furthermore, the effectiveness of the model is inherently dependent on the accuracy of pose estimation.
+Inaccurate or noisy pose predictions can introduce significant errors in the downstream recognition process.
+Nevertheless, with the rapid progress in pose estimation techniques(Huang et al., 2023; Peng et al., 2023; Zhao et al., 2023; Ai et al., 2024; Yang et al., 2023; Qiu et al., 2023; Liao et al., 2024; Tang et al., 2023), the impact of such errors has been substantially mitigated in recent years.
+A.6. Future Work
+While the proposed Parsing Skeleton representation demonstrates remarkable advantages in capturing fine-grained body dynamics and achieving strong performance across diverse scenarios, there remains significant potential to further enhance its utility through advanced multimodal fusion strategies.
+In this work, we adopt a relatively straightforward fusion approach between silhouettes and Parsing Skeletons.
+However, designing a more effective and adaptive fusion mechanism to better exploit the complementary nature of these two modalities warrants deeper exploration.
+Moreover, to emphasize the generalization and plug-and-play nature of the Parsing Skeleton, we employ two widely adopted baseline gait recognition models as the downstream architecture.
+To fully unlock the potential of this representation, future efforts could focus on designing specialized network architectures tailored to the structural and semantic characteristics of Parsing Skeletons.
+Additionally, Parsing Skeleton inherently encodes subtle motion information such as limb orientation, speed, and localized dynamics, which are crucial for fine-grained gait analysis.
+Future research could further explore these temporal and spatial cues to capture micro-movements and enhance the temporal modeling of gait sequences.